# Integrated Sensing and Communication for Vehicular Networks: A Rate-Distortion Fundamental Limits of State Estimator thanks:

### v1

@@ Abstract
The state-dependent memoryless channel (SDMC) is employed to model the integrated sensing and communication (ISAC) system for connected vehicular networks, where the transmitter conveys messages to the receiver while simultaneously estimating the state parameter of interest via the received echo signals. However, the performance of sensing has often been neglected in existing works. To address this gap, we establish the rate-distortion function for sensing performance in the SDMC model, which is defined based on standard information-theoretic principles to ensure clear operational meaning. In addition, we propose a modified Blahut-Arimoto type algorithm for solving the rate-distortion function and provide convergence proofs for the algorithm. We further define the capacity-rate-distortion tradeoff region, which, for the first time, unifies information-theoretic results for communication and sensing within a single optimization framework. Finally, we numerically evaluate the capacity-rate-distortion region and demonstrate the benefit of coding in terms of estimation rate for certain channels.
@@ I Introduction
Future vehicular networks are poised to integrate sensing and communication functionalities into unified systems [1], [2], [3]. For instance, autonomous vehicles acquire real-time environmental data through integrated networks, which facilitates navigation and helps avoid traffic congestion [4]. This paradigm, termed Integrated Sensing and Communication (ISAC), leverages shared spectrum and hardware so that the same radio signals are used both to transmit data and to sense the environment, thereby enabling simultaneous message delivery and receiver-state estimation.
@@
In the context of ISAC systems, we aim to determine, in the sense of information theory, whether the waveform design that maximizes the conditional mutual information $I(\mathsf{S};\mathsf{T}|\mathsf{X})$ mentioned above is optimal. The rate-distortion theory characterizes the minimal source coding rate compatible with a given target average distortion. Indeed, Berger mentions this idea in [18, p. 9] that: “Rate-distortion theory provides knowledge about how the frequency of faulty categorization will vary with the number and quality of observations”. For state sequences corresponding to the parameter of interest, the estimator is reconstructed by obtaining the noise-corrupted sequence while determining the minimum achievable rate that satisfies the target distortion. The codeword acts as side information in the estimator and acts as noise affecting the state sequence at the receiver, as illustrated in Figure 1. Previous information-theoretical work has focused on the noisy source setting without side information [19],[20],[21],[22],[23], [24]. These works imply that, in the limit of infinite blocklengths, the problem is equivalent to a conventional (noiseless) lossy source coding problem with the surrogate distortion measure. On the other hand, the source coding with side information at the source decoder in the lossy case is studied by Wyner and Ziv’s seminal work [25]. The side information at multiple receivers is further studied in [26]. More recent work studied a model in which the side information is corrupted by noise and is available only at the decoder, differing from the ISAC model [27]. In the coding theory literature, high-rate quantization and transform coding with side information available at the decoder for a noisy source is established in [28]. Nevertheless, it does not contain information-theoretic results applicable to the ISAC model.
@@ III-A Rate-Distortion Functions
where ties broken arbitrarily. And then for $\mathsf{h}_{a}^{(K)}(x^{n},t^{n})=i^{\star}$, let $\mathsf{h}_{b}^{(K)}=z_{i^{\star}}^{n}$. We obtain that $\mathsf{h}=\mathsf{h}_{a}^{(K)}\circ\mathsf{h}_{b}^{(K)}$ is the optimum estimator.
@@
where (14) folows from the nonasymptotic covering lemma [31, Lemma 5] and
@@
However, the probability on the RHS of (32) tends to $1$ and thus $\mathbb{E}[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))]>D$, thereby constradicting the fact that $\mathbb{E}\left[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))\right]\leq D$. Hence any $(n,M,D)$-estimator such that $\mathbb{E}\left[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))\right]\leq D$ should satisfies $\log K\geq I(\mathsf{T;\mathsf{Z}|\mathsf{X}})$.
@@ III-B Computation of Rate-Distortion Functions
In this subsection, we present a Blahut-Arimoto type algorithm that can be used to solve the optimization problem (6). By using the same argument in the proof of [32, Theorem 2.7.4], the mutual information $I(\mathsf{T};\mathsf{Z}|\mathsf{X})$ is a convex function of $P_{\mathsf{Z}|\mathsf{XT}}$ for fixed $P_{\mathsf{X}}$ and $P_{\mathsf{S}}$. To solve (6), we have the following optimization problem:
@@
    For fixed $Q_{\mathsf{Z}|\mathsf{X}}$, $P_{\mathsf{Z}|\mathsf{XT}}$, $F_{\mu}(P_{\mathsf{Z}|\mathsf{XT}},Q_{\mathsf{Z}|\mathsf{X}})$ is minimized by $P_{\mathsf{Z}|\mathsf{XT}}(z|xt)=P_{z|xt}(Q_{\mathsf{Z}|\mathsf{X}})$ for $(x,z,t)\in\mathcal{X}\times\mathcal{Z}\times\mathcal{T}$, where
@@
The following results show that the alternating optimization using Theorem 1 converges to the rate-distortion function that satisfies some target distortions. In the proof, we use the techiques similar to [33, 30].
@@ IV-A Binary Channels with Multiplicative Bernoulli States
where $p\min\{q,1-q\}<D\leq\min\{q,1-q\}$. This complets the proof. ∎
The numerical results of Corollary 1 are illustrated in Fig. 3 and 3 with parameters $p=0.3$ and $p=0.3$, respectively. The randomized estimator achieves the boundary of $\mathcal{CRD}$ by using Corollary 1. We also use Algorithm 1 to obtain the numerical results. Unsurprisingly, they coincide with Corollary 1. Through figures, we find that the deterministic estimator (40), which achieves minimum distortion, is a special case of the entire tradeoff surface. In fact, when $P_{\mathsf{X}}$ or $P_{\mathsf{S}}$ is fixed, this tradeoff problem naturally degenerates into a two-dimensional tradeoff between the target distortion $D$ and the rate-distortion function $R(D)$ (see Theorem 1) or the channel capacity $C(D)$ (see [9, Section II]), respectively.
@@ IV-B General DMCs With Multiplicative States
Then, we consider a general DMC with a uniform multiplicative state. We also use the channel $\mathsf{Y}=\mathsf{S}\mathsf{X}$ and $\mathsf{T}=\mathsf{S}\mathsf{X}$, where $\mathcal{X}=\mathcal{S}=\{0,1,2,3\}$. Let $\mathsf{S}\sim P_{\mathsf{S}}$ be a discrete memoryless source with Hamming distortion. The optimal input distribution is unique. We apply [9, Theorem 4] to solve for the input distribution $P_{\mathsf{X}}$. Then, for such a $P_{\mathsf{X}}$, we use Algorithm 1 to find the optimal estimator. The numerical results of the boundary of $\mathcal{CRD}$ are illustrated in Fig. 5 and 5. Moreover, we once again observe that the randomized estimator obtained by Algorithm 1 coincides with the deterministic estimator that achieves the minimum distortion at the minimum target distortion.
@@ IV-C Real Gaussian channels with Sensing Echo Signals Reflected From The Target
where we let $k=\sqrt{3P/(M^{2}-1)}$ to satisfy the power constraint. The resulting rate-distortion curve is shown in Fig. 6. We can see that deterministic waveforms require a lower estimation rate to achieve the same expected distortion compared to 16-ary PAM waveforms at 0 dB power constranit. However, if power constranit is $10$ dB, 16-ary PAM waveforms enable the estimator to exhibit better performance when $D\leq 6$.
@@
It should be noted that in [13] and [15], the authors maximize the conditional mutual information $I(\mathsf{S};\mathsf{T}|\mathsf{X})$ to design the deterministic waveforms. For the sensing model of (52), we have $I(\mathsf{S};\mathsf{T}|\mathsf{X})=\mathbb{E}[\frac{1}{2}\log\left(1+\mathsf{X}^{2}\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}\right)]\leq\log(1+\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}\mathbb{E}[\mathsf{X}^{2}])=\log(1+\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}P)$ by the concaveity of $\log(1+x)$ and Jensen’s inequality, with equality if and only if $\mathsf{X}=\sqrt{P}$ is a constant or $P_{\mathsf{X}}(-\sqrt{P})=P_{\mathsf{X}}(\sqrt{P})=0.5$. In numerical results of Fig. 6, we use $\mathsf{X}=\sqrt{P}$. We note that at high power constraint (e.g., $10$ dB), the $16$-ary PAM waveform gives a lower estimation rate compared to the deterministic waveform under the same expected distortion constraint when $D\leq 6$. This indicates that the performance of the estimator can be improved through the coding of the transmitter, especially when compared with the waveform which maximizes $I(\mathsf{S};\mathsf{T}|\mathsf{X})$.
@@ V Conclusion and Discussion
In summary, we introduced an estimator rate-distortion function for the ISAC model, which jointly optimizes wireless communication and sensing tasks. This function, which characterizes the minimum rate required to achieve a given estimation distortion, has a clear operational meaning and provides information-theoretic insights into waveform design. In addition, we proposed an improved Blahut-Arimoto type algorithm, which is proven to converge to the rate-distortion function in this paper and is used to find the probability distribution of the optimal estimator. Finally, we establish, for the first time, a capacity-rate-distortion region that describes the achievable tradeoff between communication capacity and estimation rate, enabling us to balance communication and sensing performance in a collaborative framework.
@@ A-A Proof of Theorem 1
where the expection is with respect to $P_{\mathsf{XTZ}}=P_{\mathsf{X}}P_{\mathsf{T}|\mathsf{X}}P_{\mathsf{Z}|\mathsf{XT}}$, (56) follows from the definition of KL divergence, (57) we condition on $(\mathsf{X},\mathsf{T})$, (61) follows from the non-negativeity of KL divergence.
@@
Then, we prove (36). The mutual information $I(\mathsf{T};\mathsf{Z}|\mathsf{X})$ is a convex function of $P_{\mathsf{Z}|\mathsf{XT}}$ for fixed $P_{\mathsf{X}}$ and $P_{\mathsf{S}}$ by using the same argument in the proof of [32, Theorem 2.7.4]. We consider the Lagrange multiplier to constrain $\sum_{z\in\mathcal{Z}}P_{\mathsf{Z}|\mathsf{XT}}(z|x,t)=1$ and obtain that
@@ 2. $\bullet$
    (70) follows from expanding the conditiondl expectation $\mathbb{E}[\mathsf{d}(\mathsf{S},a)|x,t]$;

### v2

# Integrated Sensing and Communication: Rate-Distortion Fundamental Limits of State Estimator thanks:

@@ Abstract
The state-dependent memoryless channel (SDMC) is employed to model the integrated sensing and communication (ISAC) system, where the transmitter conveys messages to the receiver while simultaneously estimating the state parameter of interest via the received echo signals. However, the performance of sensing has often been neglected in existing works. To address this gap, we establish the rate-distortion function for sensing performance in the SDMC model, which is defined based on standard information-theoretic principles to ensure clear operational meaning. In addition, we propose a modified Blahut-Arimoto type algorithm for solving the rate-distortion function and provide convergence proofs for the algorithm. We further define the capacity-rate-distortion tradeoff region, which unifies information-theoretic results for communication and sensing within a single optimization framework. Finally, we numerically evaluate the capacity-rate-distortion region and demonstrate the benefit of coding in terms of estimation rate for certain channels.
@@ I Introduction
Future communication networks are poised to integrate sensing and communication functionalities into unified systems [20], [21], [27]. For instance, autonomous vehicles acquire real-time environmental data through integrated networks, which facilitates navigation and helps avoid traffic congestion [33]. This paradigm, termed Integrated Sensing and Communication (ISAC), leverages shared spectrum and hardware so that the same radio signals are used both to transmit data and to sense the environment, thereby enabling simultaneous message delivery and receiver-state estimation.
@@
In the context of ISAC systems, we aim to determine, in the sense of information theory, whether the waveform design that maximizes the conditional mutual information $I(\mathsf{S};\mathsf{T}|\mathsf{X})$ mentioned above is optimal. The rate-distortion theory characterizes the minimal source coding rate compatible with a given target average distortion. Indeed, Berger mentions this idea in [5, p. 9] that: “Rate-distortion theory provides knowledge about how the frequency of faulty categorization will vary with the number and quality of observations”. For state sequences corresponding to the parameter of interest, the estimator is reconstructed by obtaining the noise-corrupted sequence while determining the minimum achievable rate that satisfies the target distortion. The codeword acts as side information in the estimator and acts as noise affecting the state sequence at the receiver, as illustrated in Figure 1. Previous information-theoretical work has focused on the noisy source setting without side information [11],[24],[28],[3],[13], [17]. These works imply that, in the limit of infinite blocklengths, the problem is equivalent to a conventional (noiseless) lossy source coding problem with the surrogate distortion measure. On the other hand, the source coding with side information at the source decoder in the lossy case is studied by Wyner and Ziv’s seminal work [29]. The side information at multiple receivers is further studied in [15]. More recent work studied a model in which the side information is corrupted by noise and is available only at the decoder, differing from the ISAC model [14]. In the coding theory literature, high-rate quantization and transform coding with side information available at the decoder for a noisy source is established in [23]. Nevertheless, it does not contain information-theoretic results applicable to the ISAC model.
@@ III-A Rate-Distortion Functions
where ties are broken arbitrarily. And then for $\mathsf{h}_{a}^{(K)}(x^{n},t^{n})=i^{\star}$, let $\mathsf{h}_{b}^{(K)}=z_{i^{\star}}^{n}$. We obtain that $\mathsf{h}=\mathsf{h}_{a}^{(K)}\circ\mathsf{h}_{b}^{(K)}$ is the optimum estimator.
@@
where (14) folows from the nonasymptotic covering lemma [26, Lemma 5] and
@@
However, the probability on the RHS of (32) tends to $1$ and thus $\mathbb{E}[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))]>D$, thereby contradicting the fact that $\mathbb{E}\left[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))\right]\leq D$. Hence any $(n,M,D)$-estimator such that $\mathbb{E}\left[\mathsf{d}(\mathsf{S}^{n},\mathsf{h}(\mathsf{T}^{n}))\right]\leq D$ should satisfy $\log K\geq I(\mathsf{T;\mathsf{Z}|\mathsf{X}})$.
@@ III-B Computation of Rate-Distortion Functions
In this subsection, we present a Blahut-Arimoto type algorithm that can be used to solve the optimization problem (6). By using the same argument in the proof of [9, Theorem 2.7.4], the mutual information $I(\mathsf{T};\mathsf{Z}|\mathsf{X})$ is a convex function of $P_{\mathsf{Z}|\mathsf{XT}}$ for fixed $P_{\mathsf{X}}$ and $P_{\mathsf{S}}$. To solve (6), we have the following optimization problem:
@@
    For fixed $Q_{\mathsf{Z}|\mathsf{X}}$, $F_{\mu}(P_{\mathsf{Z}|\mathsf{XT}},Q_{\mathsf{Z}|\mathsf{X}})$ is minimized by $P_{\mathsf{Z}|\mathsf{XT}}(z|xt)=P_{z|xt}(Q_{\mathsf{Z}|\mathsf{X}})$ for $(x,z,t)\in\mathcal{X}\times\mathcal{Z}\times\mathcal{T}$, where
@@
The following results show that the alternating optimization using Theorem 1 converges to the rate-distortion function that satisfies some target distortions. In the proof, we use the techniques similar to [2, 10].
@@ IV-A Binary Channels with Multiplicative Bernoulli States
where $p\min\{q,1-q\}<D\leq\min\{q,1-q\}$. This completes the proof. ∎
The numerical results of Corollary 1 are illustrated in Fig. 3 and 3 with parameters $p=0.3$ and $p=0.3$, respectively. The randomized estimator achieves the boundary of $\mathcal{CRD}$ by using Corollary 1. We also use Algorithm 1 to obtain the numerical results. Unsurprisingly, they coincide with Corollary 1. Through figures, we find that the deterministic estimator (40), which achieves minimum distortion, is a special case of the entire tradeoff surface. In fact, when $P_{\mathsf{X}}$ or $P_{\mathsf{S}}$ is fixed, this tradeoff problem naturally degenerates into a two-dimensional tradeoff between the target distortion $D$ and the rate-distortion function $R(D)$ (see Theorem 1) or the channel capacity $C(D)$ (see [1, Section II]), respectively.
@@ IV-B General DMCs with Multiplicative States
Then, we consider a general DMC with a uniform multiplicative state. We also use the channel $\mathsf{Y}=\mathsf{S}\mathsf{X}$ and $\mathsf{T}=\mathsf{S}\mathsf{X}$, where $\mathcal{X}=\mathcal{S}=\{0,1,2,3\}$. Let $\mathsf{S}\sim P_{\mathsf{S}}$ be a discrete memoryless source with Hamming distortion. The optimal input distribution is unique. We apply [1, Theorem 4] to solve for the input distribution $P_{\mathsf{X}}$. Then, for such a $P_{\mathsf{X}}$, we use Algorithm 1 to find the optimal estimator. The numerical results of the boundary of $\mathcal{CRD}$ are illustrated in Fig. 5 and 5. Moreover, we once again observe that the randomized estimator obtained by Algorithm 1 coincides with the deterministic estimator that achieves the minimum distortion at the minimum target distortion.
@@ IV-C Real Gaussian Channels with Sensing Echo Signals Reflected From the Target
where we let $k=\sqrt{3P/(M^{2}-1)}$ to satisfy the power constraint. The resulting rate-distortion curve is shown in Fig. 6. We can see that deterministic waveforms require a lower estimation rate to achieve the same expected distortion compared to 16-ary PAM waveforms at 0 dB power constraint. However, if power constraint is $10$ dB, 16-ary PAM waveforms enable the estimator to exhibit better performance when $D\leq 6$.
@@
It should be noted that in [4] and [31], the authors maximize the conditional mutual information $I(\mathsf{S};\mathsf{T}|\mathsf{X})$ to design the deterministic waveforms. For the sensing model of (52), we have $I(\mathsf{S};\mathsf{T}|\mathsf{X})=\mathbb{E}[\frac{1}{2}\log\left(1+\mathsf{X}^{2}\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}\right)]\leq\log(1+\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}\mathbb{E}[\mathsf{X}^{2}])=\log(1+\sigma_{\mathsf{S}}^{2}/\sigma_{\mathsf{N}}^{2}P)$ by the concavity of $\log(1+x)$ and Jensen’s inequality, with equality if and only if $\mathsf{X}=\sqrt{P}$ is a constant or $P_{\mathsf{X}}(-\sqrt{P})=P_{\mathsf{X}}(\sqrt{P})=0.5$. In numerical results of Fig. 6, we use $\mathsf{X}=\sqrt{P}$. We note that at high power constraint (e.g., $10$ dB), the $16$-ary PAM waveform gives a lower estimation rate compared to the deterministic waveform under the same expected distortion constraint when $D\leq 6$. This indicates that the performance of the estimator can be improved through the coding of the transmitter, especially when compared with the waveform which maximizes $I(\mathsf{S};\mathsf{T}|\mathsf{X})$.
@@ V Conclusion and Discussion
In summary, we introduced an estimator rate-distortion function for the ISAC model, which jointly optimizes wireless communication and sensing tasks. This function, which characterizes the minimum rate required to achieve a given estimation distortion, has a clear operational meaning and provides information-theoretic insights into waveform design. In addition, we proposed an improved Blahut-Arimoto type algorithm, which is proven to converge to the rate-distortion function in this paper and is used to find the probability distribution of the optimal estimator. Finally, we establish a capacity-rate-distortion region that describes the achievable tradeoff between communication capacity and estimation rate, enabling us to balance communication and sensing performance in a collaborative framework.
@@ A-A Proof of Theorem 1
where the expectation is with respect to $P_{\mathsf{XTZ}}=P_{\mathsf{X}}P_{\mathsf{T}|\mathsf{X}}P_{\mathsf{Z}|\mathsf{XT}}$, (56) follows from the definition of KL divergence, (57) we condition on $(\mathsf{X},\mathsf{T})$, (61) follows from the non-negativity of KL divergence.
@@
Then, we prove (36). The mutual information $I(\mathsf{T};\mathsf{Z}|\mathsf{X})$ is a convex function of $P_{\mathsf{Z}|\mathsf{XT}}$ for fixed $P_{\mathsf{X}}$ and $P_{\mathsf{S}}$ by using the same argument in the proof of [9, Theorem 2.7.4]. We consider the Lagrange multiplier to constrain $\sum_{z\in\mathcal{Z}}P_{\mathsf{Z}|\mathsf{XT}}(z|x,t)=1$ and obtain that
@@ 2. $\bullet$
    (70) follows from expanding the conditional expectation $\mathbb{E}[\mathsf{d}(\mathsf{S},a)|x,t]$;